A group plans to start a new business that will require $240,000 for start-up capital. The individuals in the group share the cost eually. If two addiditonal people join the group, the cost per person will decrease by $4000. How many people are presently in the group?
I let:
Capital: C = 240,000
People: P
Cost per person: I = 4000 - 2P

Answer by ewatrrr(24785)


Hi
Let P represent the number in the original group
240000/P - 240000/(P+2)= $4000
240000[(P+2) - P] = 4000*P(P+2)
4000P^2 + 8000P - 480000 = 0
P^2 + 2P - 120 = 0
factoring
(P+12)(P-10) = 0
(P+12)= 0 tossing out the negative solution for count
(P-10)= 0 P = 10, original group. (Would be 12 if 2 more joined)
CHECKING our Answer***
240,000/10 = 24,000
240,000/12 = 20,000
24,000-20,000 = $4000